OpenStudy (anonymous):

line 1 contains (2, -4) and (0,2) Line 2 contains (-4,5) and (-1, 6) are the line parallel perpendicular or neuther? please help! thankss!

OpenStudy (anonymous):

Calculate the slopes using y-y1/x-x1 (pick whichever point you like to be (x,y)

OpenStudy (anonymous):

slope of line 1 = (-4-2)/(2+4) =-1 slope of line 2 = 1 equation of line 1 y+4=x-2 y=x-6 equation of line 2 y+4=-x-2 y=-x-6